OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. Write your final answer in slope-intercept form. (-3, -5) and (0, 5)

OpenStudy (anonymous):

Use the slope formula: \[m = \frac{y_{1} - y_{2}}{x_{1} - x_{2}}\] \[m = \frac{-5 - 5}{-3 - 0}\] \[m = \frac{10}{3}\] Then plug it into point slope form: \[y - y_{1} = m(x - x_{1})\] \[y - 5 = \frac{10}{3}(x - 0)\] \[y - 5 = \frac{10x}{3}\] \[y = \frac{10x}{3} + 5\] The final answer is in slope-intercept form.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

You're welcome :)